OpenStudy (anonymous):

Find all the angles between 0 and 2π that satisfy the following equation: cos(x-0.4)=3sinx I can't seem to simplify it correctly >< Help is greatly appreciated.

OpenStudy (dumbcow):

\[\cos x \cos .4 + \sin x \sin .4 = 3 \sin x\] \[\cos x \cos .4 = 3 \sin x - \sin x \sin .4\] \[\cos x \cos .4 = \sin x (3 - \sin .4)\] \[\cos .4 = \frac{\sin x}{\cos x} (3 - \sin .4)\] \[\tan x = \frac{\cos .4}{3 - \sin .4}\]

OpenStudy (anonymous):

aah thank you

OpenStudy (dumbcow):

yw :)